OpenStudy (anonymous):

A titration was performed in a lab situation. H2SO4 was titrated with NaOH. The following data was collected: mL of NaOH used = 43.2 mL concentration NaOH = 0.15 M mL H2SO4 = 20.5 mL Notice that H2SO4 releases 2 H+ per mole. What is the concentration of H2SO4? a) 0.036 M b) 0.16 M c) 0.63 M d) 6.3 M

OpenStudy (cuanchi):

1) you have to write down and balance the neutralization equation 2) calculate how many moles of NaOH you have in 43.2 mL 0.15 M. moles= Molarity x Volume (L) 3) according to the soichiometry of the reaction calculate how many moles of H2SO4 you need 4) calculate the molarity of H2SO4 (Molarity = moles/ volume in L) moles is the value that you calculate in 3) and volume you have to convert the 20.5 mL to L

OpenStudy (anonymous):

ok. thank you.

OpenStudy (cuanchi):

did you got the answer?

OpenStudy (anonymous):

no. I got the answer wrong. I thought it was 0.63 M. but that was incorrect.

OpenStudy (rushwr):

do u still need the answer?

OpenStudy (rushwr):

I think U might have got the equation wrong ! So the balanced equation for the titration would be :\[2NaOH + H _{2}SO _{4} \rightarrow Na _{2}SO _{4} + 2H _{2}O\]

OpenStudy (rushwr):

Now you find the no. of moles of NaOH as they have given us the volume and the concentrations ! Use \[C = \frac{ n }{ V }\]

OpenStudy (rushwr):

remember no of moles of NaOH is twice the number of H2SO4 ( Look at the equation) What I go for this was \[0.15 * 43.2 * 10^{-3} = 6.48 * 10^{-3}\] So now divide the no of moles of NaOH by 2 to find out the no. of moles of H2SO4 That would be 3.24 * 10 ^-3 Now again use C=n/v for H2SO4 and obtain the concentration . Always remember to convert ml to liters. So finally I obtained 0.158 which is approximately equals to 0.16 the 2nd answer !